# Progressive Growing of Video Tokenizers for Highly Compressed Latent Spaces

### v1

@@ 4.1 Reconstruction Quality
Implementation details. We train our models on our internal dataset of 300M images and 15M videos at 256$\times$256 resolution. The 4$\times$, 8$\times$, and 16$\times$ temporal compression models are trained with 17-frame training video clips sampled at 6 fps, 12 fps, and 24 fps, respectively. Since we follow MagViT-v2’s approach of using 3D Causal Convolution layers for spatiotemporal processing, videos with $1+k\text{ x }N$ frames are compressed into $1+N$ latent frames, where $k$ is the compression ratio. Following this, the 4$\times$ models compress videos from 17 video frames to 5 latent frames, similarly, 8$\times$ from 17 frames to 3, and, 16$\times$ from 17 frames to 2.
@@
  • MagViT-v2 [41]. Since the official codebase is unavailable, we implement it by ourselves and train with a setting similar to our method. We also compare against two open-sourced variants, implemented in Open-SORA [46] and Open-SORA-Plan [24]. We name these baselines OS-VAE and OSP-VAE, respectively. Similar to the original MagViT-v2 implementation, both were implemented to perform 8$\times$ spatial and 4$\times$ temporal compression. We use the pretrained models from the official codebases.
  • We also compare to CMD [43], which decomposes videos into disentangled content and motion latents. Since the weights are not publicly available, we follow the official codebase and train the models on our internal dataset.
@@
Evaluation benchmarks. We evaluate all methods for reconstruction quality on two standard video benchmarks, MCL-JCV [38] and DAVIS 2019 (Full Resolution) [6]. Both contain 30 high-resolution videos at a 16:9 aspect ratio. We select two different settings for reconstruction, one at square crops of 512$\times$512, and the other after resizing the videos to 360$\times$640. We compare the reconstruction quality on PSNR, LPIPS [45], and Fréchet video distance (FVD) [37, 11].
Comparison on base 4$\times$ compression. Before going into the realm of high compression, we evaluate all the methods on base 4$\times$ temporal compression, since most of the methods, like OS-VAE [46], OSP-VAE [24], and MagViT-v2 [41] operate at this compression ratio. From Table 1, we observe that MagViT-v2 achieves better PSNR in all cases compared to both OS-VAE and OSP-VAE. Our method, ProMAG, albeit being more efficient than MagViT-v2 (Section 3.1), achieves a similar reconstruction quality for 4$\times$ temporal compression.
@@
Effectiveness of progressive growing for boosting temporal compression. We compare our method, ProMAG, against MagViT-v2 on 2 different settings, one with 8 channels and the other with 16 channels in the latent dimension for both 8$\times$ and 16$\times$ temporal compression. From Table 1, ProMAG, both w/ and w/o GAN training for progressive growing, achieves better reconstruction quality than MagViT-v2 directly extended to high compression. It is interesting to note that the 16-channel version of ProMAG-8$\times$ can maintain a somewhat similar reconstruction quality to ProMAG-4$\times$. Qualitative analysis in Figure 6, (left) we observe that at 16$\times$ compression, MagViT-v2 is unable to preserve even the coarse structural details in the face or the dog or the human hand. Even at 8$\times$ temporal compression, in regions of very high motion like the cyclist, MagViT-v2 blends the foreground object with the background, causing noticeable motion blur in the reconstructed videos.
@@ 4.2 T2V Quality with Compressed Latent Space
Our ultimate goal is to generate a latent space that is not only compact but more importantly, good for downstream diffusion model training. To this end, we evaluate the latent spaces produced by our models by accessing the text-to-video generation quality of the Diffusion Transformer (DiT) model trained with the resulting latent spaces. In this section, we only evaluate our 4$\times$, 8$\times$, and 16$\times$ models and not the baseline MagViT-v2 at these compression ratios. We aim to investigate if DiT training quality and convergence degrades when trained with the highly compressed latent space of 16$\times$ temporal compression compared to that trained with the original latent at 4$\times$ temporal compression.
@@
Training and implementation details. Our text-to-video diffusion model is based on the standard DiT formulation [27], composed of multiple Transformer blocks, where we replace spatial self-attention with spatial-temporal self-attention blocks. We train our model on an internal dataset of images and videos. We first train DiT on 256p images for 200K iterations, then jointly train on images and videos for about 150K more iterations.
Figure 7: Text-to-Video Evaluation. Quantitative comparison of text-to-video generation model train on 4$\times$ and 16$\times$ latent space of ProMAG. (top) The graph of time taken per denoising step of the diffusion model for generating N frame video. The time is computed on H100 with Flash Attention 3 [34]. The resolution of the videos is $192\times 360$ pixels (or $24\times 40$ in latent dimension). We find that it takes the same time per diffusion step to generate a 136-frame video using the 4$\times$ latent space, as generating a 340-frame video using the 16$\times$ latent space. (bottom) Quantitative analysis different dimension on VBench [18] on generated videos with 4$\times$ and 16$\times$ compressed latents.
@@
Figure 8: Text-to-Video Generation Results. We show very vibrant videos generated by DiT on our ProMAG-16$\times$ latent space. This highlights that training DiT on highly compressed latent space can still generate videos with accurate text coherence and realistic motion.
@@
Evaluation metrics. We evaluate the quality of text-to-video generation using VBench [18]. Following the official guidelines, we generate videos with all the 946 provided text prompts and generate videos with 5 random seeds per prompt. We evaluate based on all 16 quality dimensions such as ‘dynamic_degree’, ‘background_consistency’, etc.
@@
T2V generation quality. From Figure 7(b), we find that training DiT with 16$\times$ temporally compressed latent space does not degrade the video generation quality on almost all dimensions compared to 4$\times$ compression latent. It even achieves slightly better statistics, likely due to that we can train with higher batch sizes for 16$\times$ temporally compressed latent due to a significantly lower number of tokens. Figure 8 shows two diverse and vibrant videos generated by DiT with 16$\times$ compressed latent. We find that the DiT can accurately follow the caption and generate very realistic motion.
@@
Efficiency. Using a more compact latent space allows DiT to operate with fewer tokens, significantly reducing GPU memory and processing time. In a fixed frame length scenario (68 frames,  2.8 seconds at 24 fps), DiT with 4$\times$ compression generates 20 latents at 0.99s per timestep, while 8$\times$ compression reduces this to 12 latents at 0.58s (1.7× speedup). With 16$\times$ compression, only 8 latents are needed, achieving 0.39s per timestep (2.5× speedup). In a fixed token budget of 40 latent frames, the 4$\times$ model generates a 136-frame video ( 5.6s), whereas the 16$\times$ model produces 340 frames ( 14.1s), a 2.5× increase in video length.
@@ 5 Discussion
Table 2: Progressive Growing v/s Progressive training. Comparison of reconstruction quality of our full method against the naive way of progressive training on the 512x512 patches of the MCL-JCV video dataset.
We want to investigate if the effectiveness of our method comes from our specific design towards reusing pretrained high-quality lower compression model and only learning the remaining information in the bottleneck layers, or simply just from progressive training of the model in stages by adding more bottleneck layers (w/o skip information flow).
Progressive training. In this case, we start from a pretrained 4$\times$ model, and then add bottleneck downsampling and upsampling blocks to the encoder and decoder respectively to grow the compression rate and only train the additional blocks. This does not include the use of subsampled frames encoding or AdaNorm as our full method (Figure 4).
@@
Figure 9: Qualitative Comparison with External Interpolation. Reconstruction comparison of our method ProMAG with 16$\times$ temporal compression on a 24fps video, against a baseline where we first encode the video at low fps (with frame subsampling $s_{f}$), in this case transforming the 24fps video into 6fps with $s_{f}=4$, followed by using external interpolation method to generate the 4 in-between frames. (middle) The baseline with external frame interpolation produces very blurry outputs in regions of abrupt and high-intensity motion, compared to our method ProMAG-16$\times$ operating on 24fps video directly (right). This implies in some sense, that our progressive growing approach has in some sense learned a better interpolation for reconstruction.
@@
Table 3: Quantitaive Comparison with External Interpolation. Reconstruction quality on MCL-JCV dataset on 512x512 crops on two different settings, one at 8$\times$ and the other at 16$\times$ temporal compression.
@@
Evaluation. From Table 3, we see that our method for 8$\times$ temporal compression achieves much better reconstruction quality on PSNR, LPIPS, and FVD compared to using our 4$\times$ model on subsampled frames followed by 2X interpolation. A similar trend can be observed in the case of 16$\times$ temporal compression.
@@ 6 Conclusion and Limitations
In this work, we push the boundary to which we can perform temporal compression while keeping latent channel dimension constant. To this end, we propose a novel method of progressively growing our 4$\times$ temporal compression video tokenizer, ProMAG, to 8$\times$ and subsequently 16$\times$ temporal compression. Through extensive evaluation of reconstruction and text-to-video generation, we show that (i) Our video tokenizer can perform much better reconstruction than baseline methods at very high temporal compression rates, (ii) our high compression latent space is suitable for DiT training and provides an immense boost in terms of efficiency for video generation. However, there still exists limitations, which offers opportunities for future research. In scenarios of very high and abrupt motion, our high-compression tokenizers, while perform significantly better than the directly trained baselines, also suffers from temporal artifacts in reconstruction. In addition, in DiT we find that training on increasingly more compressed latent space tends to slightly increase the time needed for convergence.
@@ Appendix A Progressive Growing - from 8$\times$ to 16$\times$ Temporal Compression
In Figure 2 we illustrate our method of growing the temporal compression of the 4$\times$ video tokenizer to 8$\times$ temporal compression. In Figure 11, we show the details of our entire method, i.e., growing the base 4$\times$ model to 8$\times$ then ultimately to 16$\times$ temporal compression in 3 stages. Note that all the parameter weights in the subsequent are initialized from their corresponding parameters in the previous stage. The newly added parameters are initialized with random weights.
@@ Appendix B ProMAG Implementation Details
Model Architecture. In Table 4, we describe the details of our tokenizer architecture for 4$\times$, 8$\times$ and 16$\times$ temporal compression models. The encoder and decoder design of our model, ProMAG follows MagViT-v2 structure. The discriminator is taken from Stable Diffusion VAE [32], where we replace the Conv2D with Conv3D.
@@
  • Video input: 17 frames, frame stride 1, $256\times 256$ resolution.
@@ Appendix C Text-to-Video Implementation Details
Model Architecture. Our text-to-video the diffusion model is based on the standard DiT [27], composed of multiple Transformer blocks, where we replace spatial self-attention with spatial-temporal self- attention blocks. The model architecture is similar to the diffusion transformer in CogVideoX [40]. Following them, we also keep the number of model parameters to be around 5B.
@@
Training on Longer Videos. In Section 4.2, we discuss that due to the highly compact nature of our 16$\times$ latent space, we can use the same token budget for 340 frames (= 20 latent frames for 16$\times$ temporal compression), which is the same as 136 frames with 4$\times$ temporal compression. To verify the generation quality of very long videos 340 frames (or 14.1s) video at 24fps, we also train a text-to-video model for this. More specifically, we only finetune our text-to-video model trained on 16$\times$ temporally compressed latent space for an additional 10K iterations on video training of 340 frames.
Appendix D Resolving Jump Issues: Overlapping Frame Reconstruction and T2V
Following MagViT-v2 [41], train ProMAG on 17 frame video, for all 4$\times$, 8$\times$ and 16$\times$ temporal compression ratio. We found that since we train the model on 17 frames, we can only do encoding and decoding with 17 (or fewer) frames. We found a similar case is true with MagViT-v2. We found that reconstruction of more than 17 frames causes deterioration in the reconstruction results beyond 17 frames, like blurring, or in some cases checkerboard-like artifacts. Thus, for reconstruction (or text-to-video generation), we need to process it in chunks of 17 frames at a time for an N-frame video. This causes a jump-like effect in regions of high-frequency details every 17 frames. This jump is much less noticeable in reconstruction results, but more noticeable in the text-to-video generation results. For video generation, this jumping effect reduces with more training iterations but persists slightly. To counteract this, we perform encoding in chunks with an overlap of a few frames (in this case of 4 frames) Figure 10. The overlapping regions in the output are blended in pixel space with linear interpolated weights. This resolves the jumping artifacts perceptually in reconstruction and video generation results.

### v2

@@ 4.1 Reconstruction Quality
Figure 6: Effectiveness of Progressive Growing at high temporal compression. Video reconstruction example for (left) 16$\times$ temporal compression, (right) 8$\times$ compression. MagViT-v2 causes loss of details and produces artifacts in the reconstructed frames, like motion blur in the dog’s and human’s hand (left) and cyclist (right). The dotted boxes show regions of the most difference.
[…]
@@
  • MagViT-v2 [47]. As the official codebase is unavailable, we implement the continuous space version ourselves and train with a setting similar to our method. This is our main baseline where we perform full-model training for 4$\times$, 8$\times$, and 16$\times$ temporal compression at $z\_dim=$ 8 and 16.
  • For fairness of comparison, we compare with SOTA tokenizers with $z\_dim=$ 8 and 16. We compare against OmniTokenizer [44], VidTok [40], Cosmos-CV (continuous version) [1], VAE in CogVideoX [46], Wan [42], Hunyuan Video [18], CV-VAE [50], and, WF-VAE [27]. All the tokenizers are continuous-space, except for OmniTokenizer, which is discrete. Additionally, all the open-sourced SOTA tokenizers operate at 4$\times$ temporal compression, except Cosmos-CV, which has both 4$\times$ and 8$\times$ temporal compression.
Method $z\_dim$ MCL-JCV DAVIS
PSNR LPIPS rFVD PSNR LPIPS rFVD
MagViT-v2 8 30.63 6.49 58.11 28.32 7.30 129.11
OmniTokenizer [44] 24.55 11.06 124.83 23.60 15.02 290.38
VidTok [40] 31.20 6.93 62.41 28.73 8.05 150.54
ProMAG 30.99 6.53 57.91 28.43 7.38 130.56
Cosmos-CV [1] 16 32.08 9.54 50.65 29.81 13.71 113.13
CogVideoX [46] 32.57 6.21 50.75 29.8 7.49 100.56
VidTok [40] 32.95 5.55 37.7 31.12 5.22 94.04
WF-VAE [27] 33.18 5.09 39.71 30.29 6.51 88.27
Wan [42] 32.27 5.15 39.33 30.15 5.29 98.49
Hunyuan [25] 33.13 4.95 36.21 30.4 5.02 97.56
CV-VAE [50] 32.83 6.31 46.35 30.07 7.93 101.94
ProMAG 32.94 4.91 34.21 30.77 4.98 93.92
Table 1: Reconstruction comparison at base 4$\times$ temporal compression. We compare our method with baselines on MCL-JCV and DAVIS under 8$\times$8$\times$4 compression on 512$\times$512 resolution.
@@
Evaluation benchmarks. We evaluate all methods for reconstruction quality on two standard video benchmarks, MCL-JCV [43] and DAVIS 2019 (Full Resolution) [7]. We select two different settings for reconstruction, one at square crops of 512$\times$512 and the other at 360$\times$640 (at native aspect ratio of 16:9). We compare the reconstruction quality on PSNR, LPIPS [49], and Fréchet video distance (FVD) [41, 12].
Method Compression $z\_dim$ MCL-JCV DAVIS
PSNR LPIPS rFVD PSNR LPIPS rFVD
MagViT-v2 8$\times$8$\times$8 8 28.47 11.17 219.36 24.02 16.7 525.1
ProMAG 30.26 8.36 96.85 26.48 11.37 360.54
MagViT-v2 8$\times$8$\times$16 8 26.15 14.54 255.51 21.82 20.54 728.04
ProMAG 28.31 11.2 183.4 23.89 16.9 608.43
Cosmos-CV [1] 8$\times$8$\times$8 16 31.05 11.79 101.36 28.27 17.87 222.61
MagViT-v2 28.51 10.62 135.57 27.18 8.84 243.19
ProMAG 32.06 6.94 63.64 28.70 8.32 194.79
MagViT-v2 8$\times$8$\times$16 16 28.51 10.62 135.57 24.17 14.84 427.52
ProMAG 30.02 9.24 115.25 25.64 13.89 320.80
Table 2: Reconstruction comparison at high temporal compression. We compare our method with baselines on MCL-JCV and DAVIS under 8$\times$8$\times$4 compression on 512$\times$512 resolution.
Effectiveness of our base 4$\times$ ProMAG. While our main goal is not to achieve the best 4$\times$ temporal compression model, we do this experiment to verify that the base 4$\times$ model we developed can serve as a strong tokenizer. From Table 1, we see that MagViT-v2 and ProMAG at $z\_dim=8$ perform better than Omnitokenizer [44] and are comparable to SOTA tokenizer VidTok [40]. Thus, we build our base model inspired from MagViT-v2 architecture. Even at $z\_dim=16$ our method performs better than SOTA methods like Cosmos-CV (continuous) [1], CogVideoX [46], and CV-VAE [50] across all metrics and datasets and comparable to very recent SOTA methods like VidTok [40] WF-VAE [27], Wan [42], and Hunyuan Video [25], which focus on creating high-quality 4$\times$ temporal compression video tokenizers. Thus, our base 4$\times$ tokenizer is a competitive model in terms of reconstruction quality.
@@
Effectiveness of progressive growing for boosting temporal compression. We compare our method, ProMAG, against MagViT-v2 on 2 different settings, one with 8 channels and the other with 16 channels in the latent dimension for both 8$\times$ and 16$\times$ temporal compression. From Table 2, ProMAG achieves better reconstruction quality than MagViT-v2 directly extended to high compression. It is interesting to note that the 16-channel version of ProMAG-8$\times$ can maintain a somewhat similar reconstruction quality to ProMAG-4$\times$. Qualitative analysis in Figure 6, (left) we observe that at 16$\times$ compression, MagViT-v2 is unable to preserve even the coarse structural details in the face or the dog or the human hand. Even at 8$\times$ temporal compression, in regions of very high motion like the cyclist, MagViT-v2 blends the foreground object with the background, causing noticeable motion blur in the reconstructed videos. ProMAG at 8$\times$ temporal compression also performs better than Cosmos-CV.
Ablation on using GAN loss for progressive growing. In this ablation study, we verify whether adding GAN loss during progressive growing is beneficial. From Table 3, we find GAN loss at this stage while slightly improving perceptual quality (LPIPS and FVD), also slightly reduces PSNR. Additionally, not using GANs while progressive growing improves training efficiency (details in supplementary). Thus, we choose not to GAN loss during progressive growing.
Figure 7: Text-to-Video Evaluation. Quantitative comparison of T2V generation model train on 4$\times$ and 16$\times$ latent space of ProMAG. (top) The graph of time taken per denoising step of the diffusion model for generating N frame video. The time is computed on H100 with Flash Attention 3 [37]. The resolution of the videos is 192×360192$\times$ 360192 × 360 pixels. We find that it takes the same time per diffusion step to generate a 136-frame video using the 4$\times$ latent space, as generating a 340-frame video using the 16$\times$ latent space. (bottom) Quantitative analysis different dimension on VBench [19] on generated videos with 4$\times$ and 16$\times$ compressed latents.
@@ 4.2 T2V Quality with Compressed Latent Space
Method Compression $z\_dim$ MCL-JCV DAVIS
PSNR LPIPS rFVD PSNR LPIPS rFVD
ProMAG 8$\times$8$\times$8 8 32.06 6.94 63.64 28.70 8.32 194.79
ProMAG (w/ GAN) 31.78 6.21 53.18 28.21 7.76 160.22
ProMAG 8$\times$8$\times$16 16 30.02 9.24 115.25 25.64 13.89 320.80
ProMAG (w/ GAN) 29.67 8.38 99.64 25.05 12.04 274.24
Table 3: Ablation on GAN loss during progressive growing. We compare our method of progressive growing with and without using GAN loss.
@@
Our ultimate goal is to generate a latent space that is not only compact but more importantly, good for downstream diffusion model training. We evaluate the latent spaces of our tokenizers by accessing the text-to-video generation quality of the Diffusion Transformer (DiT) model trained with the resulting latent spaces. In this section, we evaluate our 4$\times$ and 16$\times$ models. We aim to investigate if DiT training quality and convergence degrades when trained with the highly compressed latent space of 16$\times$ temporal compression compared to the commonly used 4$\times$ temporal compression.
@@
Training and implementation details. Our text-to-video diffusion model is based on the standard DiT formulation [30], composed of multiple Transformer blocks, where we replace spatial self-attention with spatial-temporal self-attention blocks. We train our model on an internal dataset of images and videos. We first train DiT on 256p images for 200K iterations and then jointly train on images and videos for about 150K more iterations.
@@
Figure 8: Text-to-Video Generation Results. We show vibrant videos generated by DiT on our ProMAG-16$\times$ latent space. This highlights that training DiT on highly compressed latent space can generate videos with accurate text coherence and realistic motion.
@@
Evaluation metrics. We evaluate the quality of text-to-video generation using VBench [19]. Following the official guidelines, we generate videos with all the 946 provided text prompts and generate videos with 5 random seeds per prompt. We evaluate based on all 16 quality dimensions.
@@
T2V generation quality. From Figure 7(b), we find that training DiT with 16$\times$ temporally compressed latent space does not degrade the video generation quality on almost all dimensions compared to 4$\times$ compression latent. Figure 8 shows two diverse and vibrant videos generated by DiT with 16$\times$ compressed latent. We find that the DiT can accurately follow the caption and generate very realistic motion.
@@
Efficiency. Using a compact latent space allows DiT to operate with fewer tokens, significantly reducing GPU memory and time. In a fixed frame length case (68 frames, $\sim$2.8s at 24 fps), DiT with 4$\times$ compression generates 20 latents at 0.99s/timestep, while 8$\times$ compression reduces this to 12 latents at 0.58s (1.7$\times$ speedup). With 16$\times$, only 8 latents are needed, achieving 0.39s/timestep (2.5$\times$). In a fixed token budget of 40 latent frames, the 4$\times$ model generates a 136-frame video ( 5.6s), whereas the 16$\times$ model produces 340 frames ( 14.1s), a 2.5$\times$ increase in video length.
@@ 5 Discussion
Table 4: Progressive Growing v/s Progressive training. Comparison of reconstruction quality of our full method against the naive way of progressive training and ProMAG on the 512x512 crops of the MCL-JCV video dataset.
We want to investigate if the effectiveness of our method comes from our specific design towards reusing pretrained high-quality, lower compression model and only learning the remaining information in the bottleneck layers, or simply just from progressive training of the model in stages by adding more bottleneck layers (w/o skip information flow).
Progressive training. In this case, we start from a pretrained 4$\times$ model, and then add bottleneck downsampling and upsampling blocks to the encoder and decoder respectively, to grow the compression rate and only train the additional blocks. This does not include the use of subsampled frames encoding or AdaNorm as our full method (Figure 4). Progressive training, in essence, is similar to tokenizer in OpenSORA [51] and OpenSORA-Plan [26], which first trains a spatial-only VAE and then adds additional bottleneck 4$\times$ temporal compression blocks.
@@
Figure 9: Qualitative Comparison with External Interpolation. Reconstruction comparison of our method ProMAG with 16$\times$ temporal compression on a 24fps video, against a baseline where we first encode the video at low fps (with frame subsampling $s_{f}=4$), in this case 6fps, followed by using external interpolation method to generate the 4 in-between frames. (middle) The baseline with external frame interpolation produces very blurry outputs in regions of abrupt and high-intensity motion, compared to our method ProMAG-16$\times$ operating on 24fps video directly (right). This implies in some sense, that our progressive growing approach has in some sense learned a better interpolation for reconstruction.
@@
Table 5: Quantitaive Comparison with External Interpolation. Reconstruction quality on MCL-JCV dataset on 512x512 crops on 8$\times$ and the other at 16$\times$ temporal compression.
@@
Evaluation. From Table 5, we see that our method for 8$\times$ temporal compression achieves much better reconstruction quality compared to using our 4$\times$ model on subsampled frames followed by 2X interpolation. A similar trend can be observed in the case of 16$\times$ temporal compression.
@@ 6 Conclusion and Limitations
In this work, we push the boundary to which we can perform temporal compression while keeping latent channel dimension constant. To this end, we propose a novel method of progressively growing our 4$\times$ temporal compression video tokenizer, ProMAG, to 8$\times$ and subsequently 16$\times$ temporal compression. Through extensive evaluation of reconstruction and text-to-video generation, we show that (i) our video tokenizer can perform much better reconstruction than baseline methods at very high temporal compression rates, and (ii) our high compression latent space is suitable for DiT training and provides an immense boost in terms of efficiency for video generation. However, there still exist limitations, which offer opportunities for future research. In scenarios of very high and abrupt motion, like sports videos, our high-compression tokenizers, while performing significantly better than the directly trained baselines, also suffer from temporal artifacts in reconstruction.
@@ Appendix A Progressive Growing - from 8$\times$ to 16$\times$ Temporal Compression
In Figure 2 of the main paper we illustrate our method of growing the temporal compression of the 4$\times$ video tokenizer to 8$\times$ temporal compression. In Figure 10, we show the details of our entire method, i.e., growing the base 4$\times$ model to 8$\times$ then ultimately to 16$\times$ temporal compression in 3 stages. Note that all the parameter weights in the subsequent are initialized from their corresponding parameters in the previous stage. The newly added parameters are initialized with random weights.
@@ Appendix B ProMAG Implementation Details
Model Architecture. In Table 9, we describe the details of our tokenizer architecture for 4$\times$, 8$\times$ and 16$\times$ temporal compression models. The encoder and decoder design of our model, ProMAG follows MagViT-v2 structure. The discriminator is taken from Stable Diffusion VAE [35], where we replace the Conv2D with Conv3D. In our case, using the MagViT-v2 discriminator produced checkerboard-like artifacts in reconstunction.
@@
  • Video input: 17 frames, frame stride 1, $\times 256$ 256 resolution.
@@ Appendix C Text-to-Video Implementation Details
Model Architecture. Our text-to-video diffusion model is based on the standard DiT [30], composed of multiple Transformer blocks, where we replace spatial self-attention with spatial-temporal self- attention blocks. The model architecture is similar to the diffusion transformer in CogVideoX [46]. Following them, we also keep the number of model parameters to be around 5B.
@@
Training on Longer Videos. In Section 4.2 (Efficiency) in the main paper, we discuss that due to the highly compact nature of our 16$\times$ latent space, we can use the same token budget for 340 frames (= 20 latent frames for 16$\times$ temporal compression), which is the same as 136 frames with 4$\times$ temporal compression. To verify the generation quality of very long videos 340 frames (or 14.1s) video at 24fps, we also train text-to-video model for this. More specifically, we only finetune our text-to-video model trained on 16$\times$ temporally compressed latent space for an additional 10K iterations on video training of 340 frames.
Appendix D Resolving Jump Issues: Overlapping Frame Reconstruction and Text-to-Video Generation
Following MagViT-v2 [47], train ProMAG on 17 frame video, for all 4$\times$, 8$\times$ and 16$\times$ temporal compression ratio. We found that since we train the model on 17 frames, we can only do encoding and decoding with 17 (or less) frames. We found the similar case is true with MagViT-v2. We found that reconstruction of more than 17 frames causes deterioration in the reconstruction results beyond 17 frames, like blurring, or in some cases checkerboard-like artifacts. Thus for reconstruction (or text-to-video generation), for an N frame video, we need to process it in chunks of 17 frames at a time. This causes jumps like effect in regions of high frequency details every 17 frames. This jumps is much less noticeable in reconstruction results, but more noticeable in the text-to-video generation results. For video generation this jumping effect reduces with more training iterations but still persists slightly. To counteract this, we perform encoding in chunks with overlap of few frames (in this case of 4 frames) Figure 11. The overlapped regions in the output is blended in pixel-space with linear interpolated weights. This resolves the jumping artifacts perceptually both in reconstruction and video generation results (please refer to video results showing this in the supplementary videos).